$6,000 is put into a bank at an annual percentage rate of 13%. Find time required for investment to double if interest is compounded continously?????
I know the formula is something like A=Pe^rt where r is interest rate and t is years...but what should i do with it?

Answer by oscargut(686)

12000=6000e^(0,13t)
then
e^(0,13t)=2
then
0,13t=L(2)
t-L(2)/0,13=5,33years (aprox)